# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Abstract
Modern deep neural networks often require distributed training with many workers due to their large size. As worker numbers increase, communication overheads become the main bottleneck in data-parallel minibatch stochastic gradient methods with per-iteration gradient synchronization. Local gradient methods like Local SGD reduce communication by only syncing after several local steps. Despite understanding their convergence in i.i.d. and heterogeneous settings and knowing the importance of batch sizes for efficiency and generalization, optimal local batch sizes are difficult to determine. We introduce adaptive batch size strategies for local gradient methods that increase batch sizes adaptively to reduce minibatch gradient variance. We provide convergence guarantees under homogeneous data conditions and support our claims with image classification experiments, demonstrating the effectiveness of our strategies in training and generalization.
@@ 1 Introduction
Recent advances in deep learning, particularly large language models (LLMs) with billions of parameters, necessitate efficient distributed training on many GPUs or TPUs. Data-parallel optimization [83] simplifies this by having each worker compute and average batch gradients from different i.i.d.  data batches, followed by a global model update. Despite its efficiency with numerous workers, this method faces challenges such as costly gradient synchronization and reduced generalization from large batches. Frequent synchronization of models or gradients among workers leads to significant communication overheads, exacerbated as worker counts and model sizes increase [71, 48, 79]. This challenge intensifies with workers distributed across nodes with limited inter-node communication speeds. To mitigate this, communication-efficient distributed optimization methods [43], such as local gradient methods [67, 82, 81, 26, 73] and their variants, have been developed. In local stochastic gradient methods (Local SGD), workers train locally and synchronize less frequently by averaging models after several local steps, effectively reducing communication overheads. In data-parallel optimization, the total batch size is the sum of all workers’ local batch sizes, increasing with more workers. Large batch sizes are preferred to fully utilize worker memory, making large-batch training the standard in large-scale model training. However, this approach often results in reduced model generalization, a problem noted by LeCun et al. [47]. Large-batch training typically creates a loss landscape with sharp minima, worsening generalization [35]. Although increasing the learning rate proportionally to batch size can mitigate the generalization gap [29, 65, 66], it cannot fully close it. Keskar et al. [35] recommended adaptive sampling to gradually increase batch sizes during training [7, 19], explored in [11, 12, 44, 59]. However, this method is still confined to single- or few-worker scenarios and does not scale well with larger models and more workers.
@@
Can we develop training strategies which are communication-efficient and memory-efficient with desirable generalization performance for large-scale distributed training?
@@
We make the following contributions: (i) We develop adaptive batch size strategies for local gradient methods by extending the adaptive sampling methods in [7, 3]. Such strategies are for local batch sizes, i.e., batch sizes of individual workers. This is particularly relevant when the workers are heterogeneous devices with different computational speeds and memories, as well as different local objectives and datasets. The proposed strategies are communication-efficient by leveraging local gradient methods, and memory-efficient by leveraging large batch sizes at later stages of training. (ii) Under various standard assumptions, we provide convergence guarantees of the proposed approaches for smooth (strongly) convex and nonconvex objectives under a homogeneous data setting. (iii) We empirically demonstrate the efficacy of the proposed strategies with a three-layer convolutional neural network and ResNet for image classification on the MNIST and CIFAR-10 datasets, respectively, in terms of communication efficiency, memory efficiency, and model generalization performance. To grasp a thorough understanding of the strategies, we also study the effect of different hyperparameters such as the number of local steps and the one which controls the probability of increasing batch sizes.
@@ 2 Related Work
Minibatch SGD [13, 63, 4, 15] is the simplest and most widely used optimizer for large-scale machine learning [5], and is easily extendable with data parallelism [49]. The convergence of minibatch SGD and its variants like AdaGrad [18] and Adam [39] is well-understood across various settings, including convex [21], nonconvex [36], and nonsmooth [10] objectives, under different gradient variance assumptions. However, frequent gradient synchronization among workers leads to communication overheads, posing a significant challenge in distributed training.
@@
Local SGD [67, 83, 54], a communication-efficient variant of minibatch SGD, reduces overheads from frequent synchronization in distributed training. Recent studies [51, 17] highlight its efficacy in pre-training large language models with extensive computational resources. Optimization of Local SGD for varied data settings has shown promising convergence rates [67, 77, 76, 40, 37, 69, 74]. It includes extensions like asynchronous SGD [55, 41, 57, 45, 30] and hierarchical SGD [50, 8, 75]. Federated learning [54, 9, 33, 34, 62], another model focusing on efficiency, addresses heterogeneity and partial participation. While applicable, we defer exploring federated learning’s theoretical and empirical aspects to future work.
@@
Large-batch training not only faces communication overheads in minibatch SGD under distributed data parallelism but also shows poorer generalization despite better hardware efficiency. It has been successful in tasks like ImageNet classification and BERT training [80, 22, 64]. Addressing these, Lin et al. [50] introduced Post-local SGD, a hybrid of minibatch and Local SGD, where early training involves frequent synchronization of local models, followed by Local SGD. In large-scale environments, Ortiz et al. [58] noted that Post-local SGD does not consistently outperform Local SGD in generalization, and the transition from minibatch to Local SGD presents a trade-off between communication efficiency and generalization. Following these findings, Gu et al. [25, 24] analyzed Local SGD’s generalization via an SDE approximation and proposed the Quadratic Synchronization Rule (QSR) for adjusting local gradient steps during training.
@@
In stochastic optimization, adaptive sampling methods [7, 19, 3] adjust batch sizes based on gradient noise, enhancing training processes for smooth finite-sum problems. These methods, further explored in deep learning [11, 12, 44, 59], have not been applied to data parallelism with distributed batches. The strategies discussed here draw from batch size scheduling, a key technique in pretraining LLMs, evidenced by its use in GPT-3 [6], Nemotron-4 [60], OLMo 65B [23], and DeepSeek-V2 [14], which involve increasing batch sizes in set stages. Despite their benefits for efficiency and parallelization, these strategies are heuristic and have unclear impacts on training. This work aims to develop theoretically grounded, communication-efficient adaptive batch size strategies akin to adaptive sampling, viewed as stochastic variance-reduction methods [31, 78, 20], focusing on scalable batch size increments.
Several studies have explored adaptive batch size strategies in deep learning, including Big Batch SGD [11, 12], CABS [1], AdaBatch [16], and SimiGrad [61]. Conversely, AdaScale SGD [32] modifies learning rates for large batches instead of batch sizes. These methods, however, are often not theoretically robust, lacking solid convergence proofs or are restricted to SGD analyses under strict conditions like convexity or the Polyak–Łojasiewicz condition. Their batch size adjustment strategies usually depend on heuristic rules, such as geometric growth or decay [61], which can lead to unreliable outcomes. Crucially, these methods overlook the distributed data-parallel context, prevalent in practice, thus not addressing the needs of adaptive batch size strategies in large-scale distributed training.
@@
Besides batch sizes, recent research has also focused on optimizing learning rate schedules for Local SGD, particularly in identical data settings as discussed in [2]. This optimization is crucial in federated learning with heterogeneous data, where local objectives differ significantly in geometry and smoothness, necessitating locally adaptive learning rate schedules [56, 38]. Our adaptive local batch size strategies complement these efforts.
@@ 3.2 Minibatch and Local Stochastic Gradient Methods
Minibatch SGD under data parallelism is a fully synchronized stochastic gradient method. In particular, at each iteration $k$, after each worker $m$ has computed its local batch gradient $\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})$, the global batch gradient $\nabla F_{\mathcal{B}_{k}}(x_{k})$ is computed by averaging all local batch gradients by $\nabla F_{\mathcal{B}_{k}}(x_{k})=\frac{1}{M}\sum_{m=1}^{M}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})$ with a global batch size $b_{k}\coloneqq|\mathcal{B}_{k}|=\sum_{m=1}^{M}b_{k}^{m}=Mb_{k}^{1}$, assuming that all local batch sizes are equal. This is followed by the global model update, performed via $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$, where $\alpha_{k}$ is the learning rate of iteration $k$. The same mechanism also applies to other optimizers such as Adam [39] and AdaGrad [18].
@@
As opposed to minibatch SGD, Local SGD [67] reduces the communication frequency. Communication for model averaging (through all-reduce operations) is performed every $H$ local gradient steps, Local SGD takes the following updates:
$(\forall m\in\llbracket M\rrbracket)(\forall k\in\mathbb{N})\quad x_{k+1}^{m}=\begin{cases}\frac{1}{M}\sum_{m=1}^{M}\left(x_{k}^{m}-\alpha_{k}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\right)&\text{if }H\text{ divides }k+1,\\ x_{k}^{m}-\alpha_{k}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})&\text{otherwise}.\end{cases}$ (2)
We also consider the following alternative representation that explicitly separates the numbers of local gradient steps $H$ and communication rounds $K$. In particular, for $M$ parallel workers, $K$ rounds of communication, $H$ local gradient steps per round, Local SGD updates can also be expressed as
$(\forall m\in\llbracket M\rrbracket)(\forall k\in\llbracket 0,K-1\rrbracket)\;\begin{cases}x_{k,h+1}^{m}=x_{k,h}^{m}-\alpha_{k,h}\nabla F_{\mathcal{B}_{k,h}^{m}}(x_{k,h}^{m})&\text{for }h\in\llbracket 0,H-1\rrbracket,\\ x_{k+1,0}^{m}=\frac{1}{M}\sum_{m=1}^{M}x_{k,H}^{m}.\end{cases}$ (3)
@@ 4 Adaptive Local Batch Size Strategies for Local Gradient Methods
The proposed adaptive local batch size strategies are based on the adaptive sampling method [7] originally developed for unconstrained stochastic optimization problems using minibatch SGD. We outline its mechanism and its extension to Local SGD below.
@@ 4.1 Adaptive Sampling Methods
Introduced in [7], the adaptive sampling method for minibatch SGD was designed for a single worker ($M=1$) with the objective $F(x)=\frac{1}{n}\sum_{i=1}^{n}f(x;\xi_{i})$ over samples $\mathcal{D}=\{\xi_{i}\}_{i=1}^{n}$. This method relies on a key stochastic optimization principle: when the sample size $n$ is large, the batch gradient $\nabla F_{\mathcal{B}_{k}}$ approximates the full gradient $\nabla F$. If $f(\cdot;\xi)$ is continuously differentiable and convex for any $\xi\in\mathcal{Z}\subseteq\mathbb{R}^{p}$, then $-\nabla F_{\mathcal{B}_{k}}(x_{k})$ serves as a descent direction for $F$ at $x_{k}\in\mathbb{R}^{d}$, provided there exists $\eta\in\left(0,1\right)$ such that
@@
In practice, for each iteration $k\in\mathbb{N}$, the dynamic sample gradient algorithm [7] performs the following minibatch SGD update: $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$ with learning rate $\alpha_{k}>0$, and then checks the following (approximate) norm test condition (6), which can also be viewed as an approximation of the exact variance norm test in the stochastic setting:
@@
i.e., condition (4) holds in expectation. Here we abbreviate the conditional expectation on $\mathcal{F}_{k}$ (i.e., the filtration $\mathcal{F}_{k}\coloneqq\sigma(\{x_{0},\mathcal{B}_{0},\mathcal{B}_{1},\ldots,\mathcal{B}_{k-1}\})$) by $\mathbb{E}_{k}[\cdot]\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}]$. If condition (6) is violated, the batch size of the next iteration $b_{k+1}$ is increased, determined by the formula:
$b_{k+1}=\left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))}{\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}}\right\rceil,$
@@
Condition (7) is more relaxed than the typical assumption of uniformly bounded gradient variance for minibatch stochastic gradient methods: for any i.i.d.  batch $\mathcal{B}\subset\llbracket n\rrbracket$ and some positive constant $\sigma>0$, $\mathbb{E}_{\mathcal{B}}\left[\|\nabla F_{\mathcal{B}}(x)-\nabla F(x)\|^{2}\right]\leqslant\sigma^{2}$ for all $x\in\mathbb{R}^{d}$, which is usually unverifiable in practice. The adaptive batch size schedule using the approximate norm test may cause rapid batch size increases, undermining its purpose. To moderate these schedules, Bollapragada et al. [3] introduced the (augmented) inner product test, which manages the variance between the batch and full gradient’s inner product. Given the preference for large batch sizes for memory efficiency, controlled via $\ eta$, we focus on the norm test and defer exploring local variants of the (augmented) inner product test to future work.

### v2

@@ Abstract
Modern deep neural networks often require distributed training with many workers due to their large size. As the number of workers increases, communication overheads become the main bottleneck in data-parallel minibatch stochastic gradient methods with per-iteration gradient synchronization. Local gradient methods like Local SGD reduce communication by only synchronizing model parameters and/or gradients after several local steps. Despite an understanding of their convergence and the importance of batch sizes for training efficiency and generalization, optimal batch sizes for local gradient methods are difficult to determine. We introduce adaptive batch size strategies for local gradient methods that increase batch sizes adaptively to reduce minibatch gradient variance. We provide convergence guarantees under homogeneous data conditions and support our claims with image classification and language modeling experiments, demonstrating the effectiveness of our strategies for both training efficiency and generalization.
@@ 1 Introduction
Recent advances in deep learning, particularly large language models (LLMs) with more than billions of parameters, necessitate efficient distributed training on many workers such as GPUs or TPUs. Data-parallel optimization (Zinkevich et al., 2010) simplifies this by having each worker compute and average batch gradients from different i.i.d. data batches, followed by a global model update. Despite its efficiency with numerous workers, this method faces challenges such as costly gradient synchronization and reduced generalization from large batches. Frequent synchronization of model parameters or gradients among workers leads to significant communication overheads, exacerbated as worker counts and model sizes increase (Tang et al., 2021; Li et al., 2022; Xu et al., 2023). This challenge intensifies with workers distributed across nodes with limited inter-node communication speeds. To mitigate this, communication-efficient distributed optimization methods (Lan et al., 2020), such as local gradient methods (Stich, 2019a; Zhou and Cong, 2018; Yu et al., 2019; Haddadpour et al., 2019; Wang and Joshi, 2019) and their variants, have been developed. In local stochastic gradient methods (Local SGD), workers train locally and synchronize less frequently by averaging model parameters after several local steps, effectively reducing communication overheads. In data-parallel optimization, the total batch size is the sum of all workers’ local batch sizes, increasing with more workers. Attributed to the ever-rising memory of workers, large batch sizes are preferred to fully utilize worker memory, making large-batch training the de facto paradigm in large-scale model training. However, this approach often results in worse model generalization, a problem noted by LeCun et al. (2002). Large-batch training typically creates a loss landscape with sharp minima, worsening generalization (Keskar et al., 2017). Although scaling the learning rate with the batch size can mitigate the so-called generalization gap (Hoffer et al., 2017; Smith and Le, 2018; Smith et al., 2018), it cannot fully close it. Keskar et al. (2017) recommended adaptive sampling to gradually increase batch sizes during training (Byrd et al., 2012; Friedlander and Schmidt, 2012), explored in De et al. (2016, 2017); Lau et al. (2024); Ostroukhov et al. (2024). However, this method is still confined to single- or few-worker scenarios and does not scale well with larger models and more workers.
@@
Can we develop communication-efficient and memory-efficient training strategies with desirable generalization performance for large-scale distributed training?
@@
We make the following contributions: (i) We develop adaptive batch size strategies for local gradient methods by extending the adaptive sampling methods in Byrd et al. (2012); Bollapragada et al. (2018). Such strategies are for local batch sizes, i.e., batch sizes of individual workers. This is particularly relevant when the workers are heterogeneous devices with different computational speeds and memories, as well as different local objectives and datasets. The proposed strategies are communication-efficient by leveraging local gradient methods, and memory-efficient by leveraging large batch sizes at later stages of training. (ii) Under various standard assumptions, we provide convergence guarantees of the proposed approaches for smooth (strongly) convex and nonconvex objectives under a homogeneous data setting. (iii) We empirically demonstrate the efficacy of the proposed strategies with ResNet for image classification on the CIFAR-10 and ImageNet datasets and MicroLlama on the C4 dataset respectively, in terms of communication efficiency, memory efficiency, and model generalization performance. To grasp a thorough understanding of the strategies, we also study the effect of different hyperparameters such as the number of local steps and the one which controls the probability of increasing batch sizes.
@@ 2 Related Work
Minibatch SGD (Dean et al., 2012; Robbins and Monro, 1951; Bottou, 2010; Dekel et al., 2012) is the simplest and most widely used optimizer for large-scale machine learning (Bottou et al., 2018), and is easily extendable with data parallelism (Li et al., 2020). The convergence of minibatch SGD and its variants like AdaGrad (Duchi et al., 2011) and Adam (Kingma and Ba, 2015) is well-understood across various settings, including convex (Gower et al., 2019), nonconvex (Khaled and Richtárik, 2023), and nonsmooth (Davis et al., 2020) objectives, under different gradient variance assumptions. However, frequent gradient synchronization among workers leads to communication overheads, posing a significant challenge in distributed training.
@@
Local SGD (Stich, 2019a; Zinkevich et al., 2010; McMahan et al., 2017), a communication-efficient variant of minibatch SGD, reduces overheads from frequent synchronization in distributed training. Recent studies (Liu et al., 2024; Douillard et al., 2023) highlight its efficacy in pre-training large language models with extensive computational resources. Optimization of Local SGD for varied data settings has shown promising convergence rates (Stich, 2019a; Woodworth et al., 2020b, a; Koloskova et al., 2020; Khaled et al., 2020; Stich and Karimireddy, 2020; Wang and Joshi, 2021). It includes extensions like asynchronous SGD (Mishchenko et al., 2022; Koloskova et al., 2022; Nguyen et al., 2022; Leconte et al., 2024; Islamov et al., 2024) and hierarchical SGD (Lin et al., 2020; Castiglia et al., 2021; Wang et al., 2022). Federated learning (McMahan et al., 2017; Chen et al., 2021; Kairouz et al., 2021; Karimireddy et al., 2020; Reddi et al., 2021), another model focusing on efficiency, addresses heterogeneity and partial participation. While applicable, we defer exploring federated learning’s theoretical and empirical aspects to future work.
@@
Large-batch training not only faces communication overheads in minibatch SGD under distributed data parallelism but also shows poorer generalization despite better hardware efficiency. It has been successful in tasks like ImageNet classification and language model training (You et al., 2020; Goyal et al., 2017; Akiba et al., 2017; Shallue et al., 2019; Liu et al., 2019; Puri et al., 2018). Addressing these, Lin et al. (2020) introduced Post-local SGD, a hybrid of minibatch and Local SGD, where early training involves frequent synchronization of local models, followed by Local SGD. In large-scale environments, Ortiz et al. (2021) noted that Post-local SGD does not consistently outperform Local SGD in generalization, and the transition from minibatch to Local SGD presents a trade-off between communication efficiency and generalization. Following these findings, Gu et al. (2024, 2023) analyzed Local SGD’s generalization via an SDE approximation and proposed the Quadratic Synchronization Rule (QSR) for adjusting local gradient steps during training.
@@
In stochastic optimization, adaptive sampling methods (Byrd et al., 2012; Friedlander and Schmidt, 2012; Bollapragada et al., 2018) adjust batch sizes based on gradient noise, enhancing training processes for smooth finite-sum problems. These methods, further explored in deep learning (De et al., 2016, 2017; Lau et al., 2024; Ostroukhov et al., 2024), have not been applied to data parallelism with distributed batches. The strategies discussed here draw from batch size scheduling, a key technique in pretraining LLMs, evidenced by its use in GPT-3 (Brown et al., 2020), Nemotron-4 (Parmar et al., 2024), OLMo 65B (Groeneveld et al., 2024), and DeepSeek-V2 (DeepSeek-AI, 2024), which involve increasing batch sizes in set stages. Despite their benefits for efficiency and parallelization, these strategies are heuristic and have unclear impacts on training. This work aims to develop theoretically grounded, communication-efficient adaptive batch size strategies akin to adaptive sampling, viewed as stochastic variance-reduction methods (Johnson and Zhang, 2013; Xiao and Zhang, 2014; Gower et al., 2020), focusing on scalable batch size increments.
Several studies have explored adaptive batch size strategies in deep learning, including Big Batch SGD (De et al., 2016, 2017), CABS (Balles et al., 2017), AdaBatch (Devarakonda et al., 2017), and SimiGrad (Qin et al., 2021). Conversely, AdaScale SGD (Johnson et al., 2020) modifies learning rates for large batches instead of batch sizes. These methods, however, are often not theoretically robust, lacking solid convergence proofs or are restricted to SGD analyses under strict conditions like convexity or the Polyak–Łojasiewicz condition. Their batch size adjustment strategies usually depend on heuristic rules, such as geometric growth or decay (Qin et al., 2021), which can lead to unreliable outcomes due to the lack of theoretical guarantees. Crucially, these methods overlook the distributed data-parallel context, prevalent in practice, thus failing to address the needs of adaptive batch size strategies in large-scale distributed training.
@@
Besides batch sizes, recent research has also focused on optimizing learning rate schedules for Local SGD, particularly in identical data settings as discussed in Balles et al. (2024). This optimization is crucial in federated learning with heterogeneous data, where local objectives differ significantly in geometry and smoothness, necessitating locally adaptive learning rate schedules (Mukherjee et al., 2024; Kim et al., 2024). Our adaptive local batch size strategies complement these efforts.
@@ 3.2 Minibatch and Local Stochastic Gradient Methods
Minibatch SGD under data parallelism is a fully synchronized stochastic gradient method. In particular, at each iteration $k$, after each worker $m$ has computed its local batch gradient $\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})$, the global batch gradient $\nabla F_{\mathcal{B}_{k}}(x_{k})$ is computed by averaging all local batch gradients by $\nabla F_{\mathcal{B}_{k}}(x_{k})=\frac{1}{M}\sum_{m=1}^{M}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k})$ with a global batch size $b_{k}\coloneqq|\mathcal{B}_{k}|=\sum_{m=1}^{M}b_{k}^{m}=Mb_{k}^{1}$, assuming that all local batch sizes are equal. This is followed by the global model update, performed via $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$, where $\alpha_{k}$ is the learning rate of iteration $k$. The same mechanism also applies to other optimizers such as Adam (Kingma and Ba, 2015) and AdaGrad (Duchi et al., 2011).
@@
As opposed to minibatch SGD, Local SGD (Stich, 2019a) reduces the communication frequency. Communication for model averaging (through all-reduce operations) is performed every $H$ local gradient steps, Local SGD takes the following updates: for $m\in\llbracket M\rrbracket)$ and $k\in\mathbb{N}$,
$x_{k+1}^{m}=\begin{cases}\frac{1}{M}\sum_{m=1}^{M}\left(x_{k}^{m}-\alpha_{k}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})\right)&\text{if }H|k+1,\\ x_{k}^{m}-\alpha_{k}\nabla F_{\mathcal{B}_{k}^{m}}(x_{k}^{m})&\text{otherwise}.\end{cases}$ (2)
We also consider the following alternative representation that explicitly separates the numbers of local gradient steps $H$ and communication rounds $K$. In particular, for $M$ parallel workers, $K$ rounds of communication, $H$ local gradient steps per round, Local SGD updates can also be expressed as: for $m\in\llbracket M\rrbracket)$ and $k\in\llbracket 0,K-1\rrbracket$,
$\begin{cases}x_{k,h+1}^{m}=x_{k,h}^{m}-\alpha_{k,h}\nabla F_{\mathcal{B}_{k,h}^{m}}(x_{k,h}^{m})&\text{for }h\in\llbracket 0,H-1\rrbracket,\\ x_{k+1,0}^{m}=\frac{1}{M}\sum_{m=1}^{M}x_{k,H}^{m}.\end{cases}$ (3)
@@ 4 Adaptive Local Batch Size Strategies for Local Gradient Methods
The proposed adaptive local batch size strategies are based on the adaptive sampling method (Byrd et al., 2012) originally developed for unconstrained stochastic optimization problems using minibatch SGD. We outline its mechanism and its extension to Local SGD below.
@@ 4.1 Adaptive Sampling Methods
Introduced in Byrd et al. (2012), the adaptive sampling method for minibatch SGD was designed for a single worker ($M=1$) with the objective $F(x)=\frac{1}{n}\sum_{i=1}^{n}f(x;\xi_{i})$ over samples $\mathcal{D}=\{\xi_{i}\}_{i=1}^{n}$. This method relies on a key stochastic optimization principle: when the sample size $n$ is large, the batch gradient $\nabla F_{\mathcal{B}_{k}}$ approximates the full gradient $\nabla F$. If $f(\cdot;\xi)$ is continuously differentiable and convex for any $\xi\in\mathcal{Z}\subseteq\mathbb{R}^{p}$, then $-\nabla F_{\mathcal{B}_{k}}(x_{k})$ serves as a descent direction for $F$ at $x_{k}\in\mathbb{R}^{d}$, provided there exists $\eta\in\left(0,1\right)$ such that
@@
In practice, for each iteration $k\in\mathbb{N}$, the dynamic sample gradient algorithm (Byrd et al., 2012) performs the following minibatch SGD update: $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$ with learning rate $\alpha_{k}>0$, and then checks the following (approximate) norm test condition (6), which can also be viewed as an approximation of the exact variance norm test in the stochastic setting:
@@
i.e., condition (4) holds in expectation. Here we abbreviate the conditional expectation on $\mathcal{F}_{k}$ (i.e., the filtration $\mathcal{F}_{k}\coloneqq\sigma(\{x_{0},\mathcal{B}_{0},\mathcal{B}_{1},\ldots,\mathcal{B}_{k-1}\})$) by $\mathbb{E}_{k}[\cdot]\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}]$. If condition (6) is violated, the batch size of the next iteration $b_{k+1}$ is increased, determined by:
$b_{k+1}=\left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))}{\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}}\right\rceil,$ (8)
@@
Condition (7) is more relaxed than the typical assumption of uniformly bounded gradient variance for minibatch stochastic gradient methods: for any i.i.d. batch $\mathcal{B}\subset\llbracket n\rrbracket$ and some positive constant $\sigma>0$, $\mathbb{E}_{\mathcal{B}}\left[\|\nabla F_{\mathcal{B}}(x)-\nabla F(x)\|^{2}\right]\leqslant\sigma^{2}$ for all $x\in\mathbb{R}^{d}$, which is usually unverifiable in practice. The adaptive batch size schedule using the approximate norm test may cause rapid batch size increases, undermining its purpose. To moderate these schedules, Bollapragada et al. (2018) introduced the (augmented) inner product test, which manages the variance between the batch and full gradient’s inner product. Given the preference for large batch sizes for memory efficiency, controlled via $\ eta$, we focus on the norm test and defer exploring local variants of the (augmented) inner product test to future work.